# LITMUS: Bayesian Lag Recovery in Reverberation Mapping with Fast Differentiable Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
Though uncertainty in the geometry and kinematics of the system mean that BLR RM requires a low-redshift anchor to measure masses, it nevertheless forms a powerful tool for measuring quasar masses to cosmological distances. Recent years have seen the close of the first generation “industrial-scale” RM surveys, namely the Sloan Digital Sky Survey (SDSS; SDSS-Shen_2015) and the Australian Dark Energy Survey (OzDES; King_2015; OZDES-DR2-Lidman_2020), with large numbers of AGN observed out to high redshifts. Such surveys measure AGN masses directly, but also use their measurements to constrain the “radius-luminosity”, or $R-L$, scaling relationship for AGN: an observed power-law between the scale of the BLR and the mean photometric luminosity of the quasar (Kaspi_2000).
@@
Though these large scale surveys collect hundreds to thousands of RM lags, they are impacted by the problem of aliasing: the emergence of multiple peaks in the lag posterior distribution that emerges from the half-year seasonal gaps in observations. Aliasing has proven to be a major issue in long-baseline RM, with significant efforts being directed towards either suppressing spurious lags in the posterior distribution (e.g. SDSS-Grier_2017; CIV_Grier_2019) or filtering out suspect lags through quality cuts (e.g. OzDES-Penton_2021; OzDES-Yu_2023; SDSS-Shen_2023; OzDES-Penton_2025). Such methods often disagree about how many sources to keep, and the most stringent of them can lead to as many as $90\%$ of sources being discarded to ensure the purity of the remaining $10\%$ (see OzDES-McDougall_2025, for a comparison and discussion).
@@
As well as extending into higher redshifts and farther sources, RM has also evolved to probe both smaller and larger physical scales of the AGN, with RM being applied to both the large dusty torus that orbits beyond the BLR (e.g. Suganuma_2006; Koshida_2014; Minezaki_2019), and also to smaller scales (roughly light-days) of the accretion disk itself (Fausnaugh_2016; OzDES-Yu_2020). RM of the accretion disk typically measures lags between observations in different photometric filters, comparing the bluer light from the hot interior of the disk to the redder outer edge, allowing both the size and temperature profile of the disk to be constrained. Where X-ray observations are available, we can also probe the emission from the black hole’s corona that is believed to be the driving signal of the AGN variability (Cackett_2007). Though RM alone only directly measures scale, it can also be used to infer properties in other dimensions. For example, recent years have used accretion disk RM to prove the viscous motion of the disk material, identifying a “long” timescale of possibly negative lag associated with the inwards migration of hot material from the disk’s cooler outer edge to its hot interior (e.g. Secunda_2023). Such measurements imply a disk that is significantly thicker than the traditional “optically thick, geometrically thin” disk assumed by the famous thin-disk model of Shakura_1973. For a review of these techniques, we direct the reader to the excellent review of RM at different scales by Cackett_2021.
@@
Since its inception, RM has seen a steadily evolving bank of tools for this time delay estimation / lag recovery task. These broadly fall into two categories: non-parametric, non-Bayesian methods that are agnostic about the underlying statistical behaviour of the light curve and parameters, and parametric methods that use our understanding of the statistical properties of the AGN’s stochastic variations to fit lags with a Bayesian generative model.
Despite its conceptual simplicity, this is far from a trivial task. Aside from the ever growing tool-chest of AGN RM, the adjacent field of estimating quasar time delays from gravitational lensing, essentially the same statistical problem, encountered enough issues to warrant the Time Delay Lens Modelling Challenge (ding_2018_TDLMC1; Ding_2021_TDLMC2), an open competition amongst a number of teams to find even a single robust and reliable method for inferring lags. Though the more rigorous parametric models have been in place for well over a decade, non-parametric methods are still often used thanks to their low numerical cost and apparent lower sensitivity to aliasing. In this paper, we show that the aliasing problem of the parametric methods is in fact overstated, and a result not of the fundamental statistical properties of RM or seasonal observations but rather a numerical artefact owing to a choice of sampling method incorrect for this problem. In this paper we provide a new code, LITMUS, which solves many of the challenges of aliasing while putting rigorous statistical constraints on its impact, discussed in detail in Section 4. We describe the statistical modelling and numerical fitting of this new method, and demonstrate this approach can both recover lags in mock data better than existing tools and that it can distinguish between true lags and spurious false positives.
@@ 2 Principles of Reverberation Mapping & Complications Therein
In the most general sense, RM is a technique of measuring the radial scale of a system, however in the context of AGN BLRs the goal is to constrain the mass of the central black hole. This is done by assuming the BLR to be thin whose radius is related to a single reverberation lag $\Delta t$, and assuming the BLR is in a virialized orbit about the SMBH such that the virial mass may be estimated by:
$M_{\rm BH}=\frac{\braket{f}c\Delta t\braket{\sigma_{v}^{2}}}{G},$ (1)
where $\braket{f}$ is a geometric / kinematic ”virial factor” calibrated with local anchors of known mass (e.g. from the $M_{\mathrm{SMBH}}-M_{\mathrm{*}}$ relationship Woo_2015; Grier_2013), $c$ is the speed of light, $G$ is the gravitational constant and $\braket{\sigma_{v}^{2}}$ is the BLRs velocity dispersion as measured from the Doppler broadening of the reverberating emission line.
Lags can be acquired directly from an AGN through RM, but can also be estimated from the AGNs monochromatic luminosity at a line-specific wavelength, $\lambda L_{\lambda}$, through the power law $R-L$ relationship: $\log_{10}{\left|\Delta t\right|}=\alpha\log_{10}{\lambda L_{\lambda}}+\beta\pm\sigma$. Accurate measurement of lags is crucial not only for inferring AGN masses, but also for constraining the slope, offset and scatter parameters ($\alpha$, $\beta$ and $\sigma$) in this relationship (e.g. OzDES-McDougall_2025).
There has been a broad spectrum of approaches in attempting to consistently and reliably measure $\Delta t$ across all AGN (see Section 3), and for LITMUS we adopt the approach of Bayesian forward modelling, i.e. one in which the AGNs signals are characterised by some set of model parameters $\theta$ which we have constrained to some prior probability distribution $\pi(\theta)$, and for which any $\theta$ has a likelihood $\mathcal{L}(D|\theta)$ of reproducing our observational data $D$ such that there is a joint distribution $P(\theta|D)$ [^1] Convention differs as to how to represent represent the normalised and un-normalised joint distributions. In this paper, we use $P(\theta)$ for the normalised probability distribution, and $\mathcal{P}(\theta)$ to represent the un-normalised distribution, i.e. $P(\theta)=\frac{1}{Z}\mathcal{P}(\theta)$, where $Z$ is the model evidence per Equation 3. for parameters and observations:
@@
Methods for recovering lags within this Bayesian framework are “parametric methods”, with a few of these described in Section 3.1. This is the approach taken by LITMUS. Unique to LITMUS is the calculation of Bayesian model evidence, the total “probability mass” of the posterior distribution:
[…]
@@
As our understanding of the AGN variability has evolved we have come to describe their stochasticity as a Gaussian Process, wherein the time series observations $y$ are Gaussian-correlated observations, such that the model likelihood is:
$\mathcal{L}(\theta|D)=\frac{1}{\sqrt{(2\pi)^{N}\det(C)}}\exp(y^{T}C^{-1}y),$ (4)
where $C$ is the ‘covariance matrix’ constructed for all observations over all light curves, and $y$ is the vector of observations after subtracting off the signal means (these means also being model parameters). This includes both the intrinsic variance / covariance of the Gaussian process in the matrix $S$, and the approximately uncorrelated white noise measurement uncertainty for each observation in the diagonal matrix $N$:
$C=S+N,\;E=\delta_{ij}E_{i}E_{j}$ (5)
@@
[…]
Figure 1: A demonstration of the sort of light curves that GP modelling can reconstruct from observations. For some time-series observations (error bars) a particular GP models the entire family of underlying light curves that exhibit the power spectral density of the GP, conditioned on how well they fit the observations. In this example the light curves is fit as a DRW with $\tau=200\mathrm{d}$ and $\sigma=1$, both in arbitrary units for this demonstrative example. The shaded regions represent the 1& $1\;\&\;2\sigma$ contours of the distribution of all such walks.
@@
$\phi_{c}(t)=\sigma_{c}^{2}\exp\left(\frac{t}{\tau}\right).$ (9)
@@
Although LITMUS works only with single lags / response signals, this also extends to the multi-lag case by having response-response covariance for two response signals with transfer functions $\psi_{1}(t)$ and $\psi_{2}(t)$:
[…]
This multi-response RM sees considerable use in mapping the temperature profile of accretion disks (e.g. Yu_2020), where different colour temperatures occur at different annular radii.
@@ 2.1 The Aliasing Problem
As reverberation mapping first pressed into the regime of industrial scale RM, the endeavour soon ran afoul of the confounding effects of ‘aliasing’: a suite of problems emerging from the combination of low precision measurements (shorter exposures, looser cadence) and the half-yearly seasonal gaps in the observations. Aliasing manifests as a tendency for lag recovery methods to over-report lags at and around the “aliasing peaks”, which correspond to minimal data overlap between the continuum and response curves ($\approx 180\;\mathrm{d}$, $540\;\mathrm{d}$ etc.).
@@
The traditional approach to the aliasing problem has been to take a frequentist approach of quality control: drawing a statistical cordon around the entire lag recovery method, inclusive of the statistical model, numerical model and lag recovery software within a single ‘black box’ and determining some quality fit measure for determining how likely a recovery is to be erroneous. The OzDES team has adopted a consistent, if steadily evolving, framework for this approach: using JAVELIN as its primary lag recovery method, in conjunction with PyCCF as a validation method, and using simulation-based measures of the the likelihood of a false positive to winnow down to a high purity final data set with $<10\%$ contamination by spurious lag recoveries. Across its various data releases SDSS have trialed multiple different anti-aliasing regimes. In earlier releases (SDSS-Grier_2017; CIV_Grier_2019), they use CREAM for their lag posterior and validate through a combination of down-weighting lags that lie near aliasing peaks and rejecting recoveries that allow for non-physical negative lags. In their final data release they adopt a less stringent selection criteria, using PyROA as their primary lag recovery method and testing for significance with the $r^{2}$ recovered by PyCCF. For a full review and comparison of these selection criteria, we direct the reader to the OzDES RM wrap-up paper (OzDES-McDougall_2025).
@@
In the abstract sense, aliasing occurs because testing the half-yearly “off-season” lags coincides with little to no overlap in data between light curves. Lag recovery is as much about rejecting poor fits as identifying good ones, and such a lack of overlap means we have less chances to identify tensions that indicate such bad fits. Between seasonal gaps, over-loose interpolations are vague and weekly constraining, while over-generous interpolations are at the mercy of coincidence and can produce spurious structure. In either case, failure-modes of the light curve reconstruction / lag testing are at their worst at these difficult lags. In Figure 2 we demonstrate this for mock data, and in particular show the multimodality that occurs in the lag posterior for the specific case of a parametric Gaussian process model. Aliasing becomes worse when the true lag is near an aliasing peak, as these are more ambiguously observed and produce a shallower peak in the posterior. In such cases we can only observe tension / discontinuity in the reconstructed light curve in the joins between observation seasons, and so aliasing becomes more pronounced for more rapid timescales of variability (smaller $\tau$ in Equation 9) and for fewer observed seasons.
@@
Figure 2: A demonstration of the source of the aliasing problem, specifically in the context of a parametric GP model. Top shows mock data with cadence, measurement uncertainty and baseline similar to OzDES with a DRW timescale of $\tau=200\mathrm{d}$ and a true lag of $\Delta t=360\mathrm{d}$. From left to right the sub-panels show lags being tested at $\Delta t=0\mathrm{d}$, $180\mathrm{d}$ and $360\mathrm{d}$. The left panel is clearly a bad fit as we can near simultaneous observations being in clear tension, and the right panel is a clear good fit as we see very little tension. The middle panel, corresponding to the first aliasing peak, is an ambiguously good fit; the lack of overlap means we cannot observe clear tensions between the light curves. The bottom panel shows the (un-normalised) log-posterior distribution with all non-lag parameters fixed at their true values. At “on-season” lags (un-shaded) we can easily reject bad fits, and so the posterior is extremely low. During the off-season lags (red shading) there are local optima arising from the ambiguity. The mode associated with the true lag (red dot) is clearly defined and dominates over aliasing modes, with the rest of the posterior being $<1\%$ of the maximum posterior density in this well behaved, high SNR example. Even so, the posterior still suffers from the rough geometry and multimodality that introduces numerical challenges in navigating it.
@@
In the example shown in Figure 2, the lag is well constrained to its true value as the aliasing peaks are several orders of magnitude shallower. However some Bayesian modelling tools can fail even in these cases due to the secondary numerical challenges that aliasing poses. JAVELIN fits lags by sampling the posterior with emcee, an implementation the Affine Invariant Ensemble Sampler (Goodman_2010_AIES). Though emcee has a well earned reputation as a robust and reliable sampler, it is decidedly not suited to such multimodal distributions. Its proposal method, which involves drawing a chord between two samples and performing a “stretch-move” to propose a new sample, cannot easily mix between well separated modes. The result is the ensemble of live points becomes partially ‘pinned’ at the aliasing modes, over-sampling them and giving the false appearance of a significant bulk of posterior density, sometimes multiple orders of magnitude over the true height of the peak. In Figure 3 we see the dangerous failure mode that can result: the mirage appearance of an aliasing peak in the reported posterior where none truly exists.
@@
Figure 3: A demonstration of the failure mode of the Affine-Invariant Ensemble Sampler (AIES), the MCMC proposal algorithm used by emcee, in multi-modal distributions. Both top and bottom panels are posterior distributions generated from the same mock data with a true lag at $\Delta t=180\mathrm{d}$, with the top panel being the result from the AIES, the same MCMC sampler as JAVELIN, while the bottom is found from exhaustive sampling of the prior range. The AIES estimate for the posterior has produced an aliasing peak at $\Delta t=540\mathrm{d}$ where none truly exists due to its ensemble of live sampling points becoming pinned at this minor mode.

### v2

@@ Abstract
Reverberation mapping is a technique in which the mass of a Seyfert I galaxy’s central supermassive black hole is estimated, along with the system’s physical scale, from the timescale at which variations in brightness propagate through the galactic nucleus. This mapping allows for a long baseline of time measurements to extract spatial information beyond the angular resolution of our telescopes, and is the main means of constraining supermassive black hole masses at high redshift. The most recent generation of multi-year reverberation mapping campaigns for large numbers of active galactic nuclei (AGN) (e.g. OzDES) have had to deal with persistent complications of identifying false positives, such as those arising from aliasing due to seasonal gaps in time-series data. We introduce LITMUS (Lag Inference Through the Mixed Use of Samplers), a modern lag recovery tool built on the “damped random walk” model of quasar variability, built in the automatic differentiation framework jax. LITMUS is purpose-built to handle the multimodal aliasing of seasonal observation windows and provides Bayesian evidence integrals for model comparison and null hypothesis testing, a more quantified alternative to existing post-fit selection methods. LITMUS also offers a flexible and modular framework for using more expressive high dimensional models for the AGN variability, and includes jax-enabled implementations of other popular lag recovery methods like nested sampling and the interpolated cross correlation function. We test LITMUS on a number of mock light curves modelled after the OzDES sample and find that it recovers their lags with high precision and successfully identifies spurious lag recoveries, reducing its false positive rate to drastically outperform the state of the art program JAVELIN. LITMUS’s high performance is accomplished by an algorithm for mapping the Bayesian posterior density which both constrains the lag and provides Bayesian evidences for model comparison and null hypothesis testing while outperforming nested sampling in computational cost by an order of magnitude.
@@ 1 Introduction
Though uncertainty in the geometry and kinematics of the system mean that BLR RM requires a low-redshift anchor to measure masses, it nevertheless forms a powerful tool for measuring quasar masses to cosmological distances. Recent years have seen the close of the first generation “industrial-scale” RM surveys, namely the Sloan Digital Sky Survey (SDSS; SDSS-Shen_2015) and the Australian Dark Energy Survey (OzDES; King_2015; OZDES-DR2-Lidman_2020), with large numbers of AGN observed out to high redshifts. Such surveys measure AGN masses directly, but also use their measurements to constrain the “radius-luminosity”, or RLR-L, scaling relationship for AGN: an observed power-law between the scale of the BLR and the mean photometric luminosity of the quasar (Kaspi_2000).
@@
Though these large scale surveys collect hundreds to thousands of RM lags, they are impacted by the problem of aliasing: the emergence of multiple peaks in the lag posterior distribution that emerges from the annual six month seasonal gaps in observations. Aliasing has proven to be a major issue in long-baseline RM, with significant efforts being directed towards either suppressing spurious lags in the posterior distribution (e.g. SDSS-Grier_2017; CIV_Grier_2019) or filtering out suspect lags through quality cuts (e.g. OzDES-Penton_2021; OzDES-Yu_2023; SDSS-Shen_2023; OzDES-Penton_2025). Such methods often disagree about how many sources to keep, and the most stringent of them can lead to as many as $90\%$ of sources being discarded to ensure the purity of the remaining $10\%$ (see OzDES-McDougall_2025, for a comparison and discussion).
@@
As well as extending into higher redshifts and farther sources, RM has also evolved to probe both smaller and larger physical scales of the AGN, with RM being applied to both the large dusty torus that orbits beyond the BLR (e.g. Suganuma_2006; Koshida_2014; Minezaki_2019), and also to smaller scales (roughly light-days) of the accretion disk itself (Fausnaugh_2016; OzDES-Yu_2020). RM of the accretion disk typically measures lags between observations in different photometric filters, comparing the bluer light from the hot interior of the disk to the redder outer edge, allowing both the size and temperature profile of the disk to be constrained. Where X-ray observations are available, we can also probe the emission from the black hole’s corona that is believed to be the driving signal of the AGN variability (Cackett_2007). Though RM alone only directly measures scale, it can also be used to infer properties in other dimensions. For example, in recent years intensive broad-band RM has revealed possible evidence of viscous motion of the disk material, identifying a “long” timescale of possibly negative lag associated with the inwards migration of hot material from the disk’s cooler outer edge to its hot interior (e.g. Santisteban_2020; Secunda_2023). Such measurements imply a disk that is significantly thicker than the traditional “optically thick, geometrically thin” disk assumed by the famous thin-disk model of Shakura_1973. For a review of these techniques, we direct the reader to the excellent review of RM at different scales by Cackett_2021.
@@
Since its inception, RM has seen a steadily evolving bank of tools for this time delay estimation / lag recovery task. These broadly fall into two categories: those that model the AGN light curves as a Gaussian Process and are agnostic about the underlying statistical behaviour of the light curve and parameters, and those that use our understanding of the statistical properties of the AGN’s stochastic variations to fit lags with a Bayesian generative model.
Despite its conceptual simplicity, this is far from a trivial task. Aside from the ever growing tool-chest of AGN RM, the adjacent field of estimating quasar time delays from gravitational lensing, essentially the same statistical problem, encountered enough issues to warrant the Time Delay Lens Modelling Challenge (ding_2018_TDLMC1; Ding_2021_TDLMC2), an open competition amongst a number of teams to find even a single robust and reliable method for inferring lags. Though the more rigorous GP models have been in place for well over a decade, non-GP methods are still often used thanks to their low numerical cost and apparent lower sensitivity to aliasing. In this paper, we show that the aliasing problem of the GP methods is in fact overstated, and a result not of the fundamental statistical properties of RM or seasonal observations but rather a numerical artefact owing to a choice of sampling method incorrect for this problem. In this paper we provide a new code, LITMUS, which solves many of the challenges of aliasing while putting rigorous statistical constraints on its impact, discussed in detail in Section 4. We describe the statistical modelling and numerical fitting of this new method, and demonstrate that this approach can both recover lags in mock data better than existing tools and that it can distinguish between true lags and spurious false positives.
@@ 2 Principles of Reverberation Mapping & Complications Therein
In the most general sense, RM is a technique of measuring the radial scale of a system, however in the context of surveys like OzDES the goal is to constrain the mass of the central black hole via the size of the BLR. This is done by simplifying the BLR to be a thin disk with a single characteristic radius/ reverberation lag $R=c\cdot\Delta t$, where $c$ is the speed of light. Assuming the BLR is in a virialized orbit about the SMBH such, the virial mass can be estimated with a line-of-sight velocity dispersion from the doppler broadening of the emission line profile. The unknown geometry and kinematics are captured in the “virial factor”, calibrated with local anchors of known mass (e.g. from the $M_{\mathrm{SMBH}}-M_{\mathrm{*}}$ relationship; Woo_2015; Grier_2013). This factor has high population dispersion, imposing on RM mass measurements an uncertainty floor of roughly a quarter. RM lags are also used to constrain the power-law scaling relationship between BLR radius and AGN luminosity (the RLR-L relationship). Accurate measurement of lags is crucial not only for inferring AGN masses, but also for constraining the slope, offset and scatter parameters in this relationship (e.g. OzDES-McDougall_2025).
There has been a broad spectrum of approaches in attempting to consistently and reliably measure $\Delta t$ across all AGN (see Section 3), and for LITMUS we adopt the approach of Bayesian forward modelling, i.e. one in which the AGNs signals are characterised by some set of model parameters $\theta$ which we have constrained to some prior probability distribution $\pi(\theta)$, and for which any $\theta$ has a likelihood $\mathcal{L}(D|\theta)$ of reproducing our observational data $D$ such that there is a joint distribution $P(\theta|D)$ [^1] In this paper, we use $P(\theta)$ for the normalised probability distribution, and $\mathcal{P}(\theta)$ to represent the un-normalised joint distribution, i.e. $P(\theta)=\frac{1}{Z}\mathcal{P}(\theta|)$, where $Z$ is the model evidence per Equation 2. for parameters and observations:
@@
We describe a few existing methods for recovering lags within this Bayesian GP framework in Section 3.1. This is the approach taken by LITMUS. Unique to LITMUS is the calculation of Bayesian model evidence, the total “probability mass” of the posterior distribution:
$Z=\int\mathcal{P}(\theta|D)d\theta.$ (2)
@@
As our understanding of the AGN variability has evolved we have come to describe their stochasticity as a GP, wherein the time series observations $y$ are Gaussian-correlated observations, such that the model likelihood is:
$\mathcal{L}(\theta|D)=\frac{1}{\sqrt{(2\pi)^{N}\det(C)}}\exp\left(-\frac{1}{2}y^{T}C^{-1}y\right),$ (3)
where $C$ is the ‘covariance matrix’ constructed for all observations over all light curves, and $y$ is the vector of observations after subtracting off the signal means (these means also being model parameters). This includes both the intrinsic variance / covariance of the GP in the matrix $S$, and the approximately uncorrelated white noise measurement uncertainty for each observation ($E_{i}$ for measurement $i$) in the diagonal matrix $N$:
$C=S+N,\;N=\delta_{ij}E_{i}E_{j}$ (4)
@@
$\phi_{rc}(t)=\iint_{-\infty}^{\infty}{\phi_{c}(t)\psi(t^{\prime})\psi(t^{\prime\prime})}dt^{\prime}dt^{\prime\prime}.$ (7)
Figure 1: A demonstration of the sort of light curves that GP modelling can reconstruct from observations. For some time-series observations (error bars) a particular GP models the entire family of underlying light curves that exhibit the power spectral density of the GP, conditioned on how well they fit the observations. In this example the light curves is fit as a DRW with $\tau=200\mathrm{d}$ and $\sigma=1$, both in arbitrary units for this demonstrative example. The shaded regions represent the 11 and $2\sigma$ contours of the distribution of all such walks.
@@
$\phi_{c}(t)=\sigma_{c}^{2}\exp\left(\frac{-\lvert t\rvert}{\tau}\right).$ (8)
@@
Although LITMUS works only with single lags / response signals in its initial release, this model also extends to the multi-lag case by having response-response covariance for two response signals with transfer functions $\psi_{1}(t)$ and $\psi_{2}(t)$:
$\phi_{rr}(t)=\iint_{-\infty}^{\infty}{\phi(t)\psi_{1}(t^{\prime})\psi_{2}(t^{\prime\prime})}dt^{\prime}dt^{\prime\prime}.$ (9)
This multi-lag RM sees considerable use in mapping the temperature profile of accretion disks (e.g. Yu_2020), where different colour temperatures occur at different annular radii.
@@ 2.1 The Aliasing Problem
As reverberation mapping first pressed into the regime of industrial scale RM, the endeavour soon ran afoul of the confounding effects of ‘aliasing’: a suite of problems emerging from the combination of low precision measurements (shorter exposures, looser cadence) and the half-yearly seasonal gaps in the observations. Aliasing manifests as a tendency for lag recovery methods to over-report lags at and around the “aliasing peaks”, which correspond to minimal data overlap between the continuum and response curves ($\approx 180\;\mathrm{d}$, $540\;\mathrm{d}$ etc.). This aliasing problem is a major confounding factor in modern RM, with the arising false positive lags dominating the mass measurements and RLR-L constraints when false positives are not screened.
@@
The traditional approach to the aliasing problem has been to take a frequentist approach of quality control: drawing a statistical cordon around the entire lag recovery method, inclusive of the statistical model, numerical model and lag recovery software within a single ‘black box’ and determining some fit quality measure for determining how likely a recovery is to be erroneous. The OzDES team has adopted a consistent, if steadily evolving, framework for this approach: using JAVELIN as its primary lag recovery method, in conjunction with PyCCF as a validation method, and using simulation-based measures of the the likelihood of a false positive to winnow down to a high purity final data set with $10\%$ contamination by spurious lag recoveries. Across its various data releases SDSS have trialled multiple different anti-aliasing regimes. In earlier releases (SDSS-Grier_2017; CIV_Grier_2019), they use CREAM (Starkey_2015_CREAM) for their lag posterior and validate through a combination of down-weighting lags that lie near aliasing peaks and rejecting recoveries that allow for non-physical negative lags. In their final data release they adopt a less stringent selection criteria, using PyROA (pyroa-ferus_2021) as their primary lag recovery method and testing for significance with the $r^{2}$ recovered by PyCCF. For a full review and comparison of these selection criteria, we direct the reader to the OzDES RM wrap-up paper (OzDES-McDougall_2025).
@@
In the abstract sense, aliasing occurs because testing the half-yearly “off-season” lags coincides with little to no overlap in data between light curves. Lag recovery is as much about rejecting poor fits as identifying good ones, and such a lack of overlap means we have fewer chances to identify tensions that indicate such bad fits. Between seasonal gaps, over-loose interpolations are vague and weekly constraining, while over-generous interpolations are at the mercy of coincidence and can produce spurious structure. In either case, failure-modes of the light curve reconstruction / lag testing are at their worst at these difficult lags. In Figure 2 we demonstrate this for mock data, and in particular show the multimodality that occurs in the lag posterior for the specific case of a parametric GP model. Aliasing becomes worse when the true lag is near an aliasing peak, as these are more ambiguously observed and produce a shallower peak in the posterior. In such cases we can only observe tension / discontinuity in the reconstructed light curve in the joins between observation seasons, and so aliasing becomes more pronounced for more rapid timescales of variability (smaller $\tau$ in Equation 8) and for fewer observed seasons.
@@
Figure 2: A demonstration of the source of the aliasing problem, specifically in the context of a parametric GP model. Top shows mock data with cadence, measurement uncertainty and baseline similar to OzDES with a DRW timescale of $\tau=200\mathrm{d}$ and a true lag of $\Delta t=360\mathrm{d}$. From left to right the sub-panels show lags being tested at $\Delta t=0\mathrm{d}$, $180\mathrm{d}$ and $360\mathrm{d}$. The left panel is clearly a bad fit as near simultaneous observations are in clear tension, and the right panel is a clear good fit as we see very little tension. The middle panel, corresponding to the first aliasing peak, is an ambiguously good fit; the lack of overlap means we cannot observe clear tensions between the light curves. The bottom panel shows the (un-normalised) log-natural of the posterior distribution, with all non-lag parameters fixed at their true values. At “on-season” lags (un-shaded) we can easily reject bad fits, and so the posterior is extremely low. During the off-season lags (blue shading) there are local optima arising from the ambiguity. The mode associated with the true lag (red dot) is clearly defined and dominates over aliasing modes, with the rest of the posterior being $1\%$ of the maximum posterior density in this well behaved, high SNR example. Even so, the posterior still suffers from the rough geometry and multimodality that introduces numerical challenges in navigating it.
@@
In the example shown in Figure 2, the lag is well constrained to its true value as the aliasing peaks are several orders of magnitude shallower. However some Bayesian modelling tools can fail even in these cases due to the secondary numerical challenges that aliasing poses. JAVELIN fits lags by sampling the posterior with emcee, an implementation of the Affine Invariant Ensemble Sampler (Goodman_2010_AIES). Though emcee has a well earned reputation as a robust and reliable sampler, it is decidedly not suited to such multimodal distributions. Its proposal method, which involves drawing a chord between two samples and performing a “stretch-move” to propose a new sample, cannot easily mix between well separated modes. The result is the ensemble of live points becomes partially ‘pinned’ at the aliasing modes, over-sampling them and giving the false appearance of a significant bulk of posterior density, sometimes multiple orders of magnitude over the true height of the peak. In Figure 3 we see the dangerous failure mode that can result: the mirage appearance of an aliasing peak in the reported posterior where none truly exists.
@@
Figure 3: A demonstration of the failure mode of the Affine-Invariant Ensemble Sampler (AIES), the MCMC proposal algorithm used by emcee, in multi-modal distributions. Both top and bottom panels are posterior distributions generated from the same mock data with a true lag at $\Delta t=854\mathrm{d}$ (dashed line), with the bottom panel being the result from the AIES, the same MCMC sampler as JAVELIN, while the top is found from exhaustive sampling of the prior range. The AIES estimate for the posterior has produced an aliasing peak at $\Delta t=540\mathrm{d}$ where none truly exists due to its ensemble of live sampling points becoming pinned at this minor mode.